Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Intern. region
Vertex Tracker
Drift chamber
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Friday April 9 Background Remediation video Meeting

1/ Report form the first BABAR/PEP-II common meeting held Wednesday April 7

John Seeman reported that the model he presented was agreed to:

-Turn-on PEP-II while minimising radiation damage , not optimising machine performances

-Stay at very low currents (10 mA) for all the debugging period

-Turn the solenoid on in stages (first level 1/3)

The action Items from the meeting were discussed:

1-Ensure availability of full background sensors and monitoring

for machine studies shifts. (Mattison, Wormser). A table was asked by the machine people.

The latest information can be found here and is maintained by T. Geld and A. McKemey

2-Plan for having no access to detector after 5 May except for major

problems. (Smith) [Also for restricted access from 30 April - 5 May]

3-Reschedule BaBar operations meetings to avoid conflict with

0830 daily fora. (Walt)

The idea is to have the 8:00 MCC meeting running in parallel with teh Bbar operation meeting and then a common meeting to decide the planning of the day at 9;00.

We need to prepare a summary report sheet for the MCC 8:00 meeting.

4-Prepare first-draft radiation budgets for next meeting (Tues April 13)


John Seeman indicated his proposed guidelines to determine the threshold levels: 10 years of operations, 30 weeks a year, threshold set to 3 times the average maximum, comparison made each shift. GW to discuss with SVT and EMC groups to prepare Tuesday meeting.

2/ Training week aftermath

1/ Video available, talks soon available on the web

There will no formal training week First week of may since most people (20 out of 27) were there at least partially during the first week.

2/ Liaison home page and hypernews group available thanks to Terry Geld and Tim Meyer.Please subscribe, bfmail list on the way

3/ Tom Mattison will send a message around asking for personal schedule, veto etc. Please answer promptly. He indicated that shift liaisons may not start exactly when the machine starts..

4/ Action items for issues raised during the week

  1. Policy wrt E150 and E157 (BABAR management)

D. Hitlin said he would investigate. It was reported during the meeting that BABAR would have 1st priority.

2. Do we need a "pep-ii tuning" state ,to which babar would react by

turning off SVT. Can we get by with a 2 threshold system , a low one when

SVT is on and a higher one otherwise (T. Mattison)

The conclusion is that there is no need for extra software. The tuning state will be a ‘injection request’ by which BABAR has to react by loading the proper config file (normal injection files if we have to refill from scratch or special tuning file if the fill is there). The two-threshold system is already available. It will be triggered by the SVT bias command, rather than readback to start with. The usefulness of the two thresholds remains to be clearly demonstrated.

  1. BABAR Xray detection strategy (P. Burchat, C. Hast, A. Roodman)

There is a need for MC studies to look for specific patterns caused by X rays in each system. PIN diodes system look promising. Technically, Mike Sullivan has to generate a proper X ray incoming rays and Terry Geld has to discuss with him an easy interface with BBSIM.

Witold pointed out that somemachine experiments like tuning x or x’ at the IP could change significantly the X ray rate so it could be very useful. Witold to check with Mike Sullivan

4. Liaison account (Tom Pavel)

Tom Pavel answered that a babar account similar to the one existing in IR2 will be created for us.

5. Copy of run control screen)(for the liaison to know how babar-daq is

doing) (N. Geddes)

Neil said it was doable and will be available.

6. Alarm handler with simplifed config file (J. Olsen)

Jim agreed to do that, provided Terry gives him the list of variables names.

7. Damage assessment (SVT and EMC groups)

the following messages were received from the SVT and EMC groups. Work is clearly progressing in this area.

8. Injection interlock using magnet green status (T. Mattison)

Tom will discuss this with PEP-II people

9. Administrative instructions to PEP-II (Q4,B3,..) (T. Mattison)



Software reports from PRV0

  1. Q2 problem (T. Geld)
  2. The Q2 was full of air because an overlap problem in GEANT volumes. It has been fixed now. No big effect was found for the calorimeter. For the DCH, introducing the material has decreased the background in the DCH end plate by a factor 2 to 3.

  3. Tracking with PRV0 (S. Wagner)
  4. The main problem due to background is the wrong bunch assignment for t0 leading to many tracks loosing hits. New algorithms are under investigation. Besides this, the loss of tracking efficiency for J/Psi K0s is 10% and the resolution is largely unaffected.

  5. Event sizes (Gregory Dubois Felsmann)

The event size was found to be 28 kBytes but with big elements missing: EMC, IFR, EMT, other parts of the trigger. The EMC will have no sparsification available at the beginning. Gregory estimates the total size to lie around 40-50 kB. This will limit the L1 trigger rate to 1 kHz.